OpenStudy (anonymous):

what is the solution to this problem ? 7/8 = n/32

OpenStudy (anonymous):

You have to cross multiply. 7/8=n/32 do the diagonals 7*32=8n then you solve the equation.

OpenStudy (anonymous):

my answer: Multiply both sides of the equation by 32. then 32 times 7 divided by 8 equals n n equals 28

OpenStudy (anonymous):

\(\bbox [ 25pt, #333 ,border: 15px solid #123 ] {\Huge\sf\color{white} {\ \ 〄EMIN∃M〄\ }}\)

OpenStudy (anonymous):

\(\bbox [ 25pt, #333 ,border: 15px solid #123 ] {\Huge\sf\color{white} {УѺҐ }Ӧ}\)

OpenStudy (anonymous):

thnx :D also it says yolo :D

OpenStudy (anonymous):

Yes you got it right.

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

my real account is @blackops2luvr i got hacked :D

OpenStudy (anonymous):

Oh well I hope you have fun on your new account ^_^

OpenStudy (anonymous):

✄☰☰☝

OpenStudy (anonymous):

\(\bbox [ 25pt, #333 ,border: 15px solid #123 ] {\Huge\sf\color{white} {❶❷❸❹❺❻❼❽❾❿ }}\)